Week of 161031

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maria D. (SCOD), Iain (Computing), Andrei (DB), Vincent (security), Luca (Storage), Andrea M. (MW), Julia (Ops), Marcelo (LHCb).
  • remote: Alexander (NL_T1), Dario (ATLAS), Kyle (OSG), Dmytro (NDGF), John K. (RAL), Sang Un Ahn (KISTI), Christoph (CMS), Di (Triumf), Fa-Hui Lin (ASGC), Victor (JINR), Xin Zhao (BNL).

Experiments round table:

  • ATLAS reports ( raw view) -
    • Quiet week. Only disruption was the kernel upgrades with the security patch. Most sites did rolling upgrades (thanks!), only a few went into complete downtimes. All in operation now.

  • CMS reports ( raw view) -
    • pp run finished last week
    • Tier-0 backlog (due to rushing with high rates in the last days) basically gone
    • Grid production went basically to zero over the weekend
      • Likely an issue with WMAgents - under investigation

  • ALICE -
    • CERN: CREAM job submission slowness and errors last Fri (GGUS:124705)

  • LHCb reports ( raw view) -
    • Activity
      • Monte Carlo simulation, data reconstruction/stripping and user jobs on the Grid
    • Site Issues
      • T0:
        • Ticket opened for Cream-CE (GGUS:123957) - in progress
      • T1:
        • CNAF: Downtime Outage wednesday and thursday LHCb SE will be unavailable.
        • NIKHEF : dCache problem is making some data unavailable (Downtime between today and tomorrow)
        • RAL : Outage tomorrow for kernel patching, Tape system will be down
        • RRCKI - very slow transfers caused by network problems, more then 100 hours data backlog (GGUS:124538) - in progress
On the last point, GGUS:124538 Marcelo mentioned that disk performance is getting better but transfers to tape are still unbearably slow. Vladimir also noticed a characteristic behavior related to VOboxes. Maria suggest to put all observations in the ticket itself to help site managers, as they didn't connect today.

Sites / Services round table:

  • ASGC: NTR
  • BNL:
    • BNL HPSS Tape system will be shutdown for software upgrade, for this week (10/31/2016, 11AM UTC to 11/04/2016, 23 PM UTC). Impact on production is minimum, as almost all data being used are staged on dCache beforehand.
    • vulnerability CVE-2016-5195, interactive nodes and compute nodes are patched via Ksplice updates
  • CNAF: (in email after the meeting) Downtime Outage wednesday and thursday LHCb SE will be unavailable. the LHCB downtime on gocdb was a mistake and has been revoked, the right one is: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=21816 it will affect ATLAS Disk, ALICE Disk and AMS Disk.
  • EGI: not connected
  • FNAL: NTR
  • GridPP: not connected
  • IN2P3: No one will connect from IN2P3 -CC today. Regarding EGI-SVG-CVE-2016-5195 : had a downtime from October 24th in the evening until 26th in the afternoon to apply patches to all services. Running without noticeable problems since then.
  • JINR: SLF6 Kernel was being upgraded "on the fly". Now finished.
  • KISTI: Scheduled downtime from 2/11 at midnight UTC to 3/11 at noon.
  • KIT:
    • Downtime on short notice for batch farm on Thursday (27.10.16) 10:00-15:00 CEST. Applied kernel patch for CVE-2016-5195 (DirtyCOW).
  • NDGF:
    • Reboot of some storage machines due to COW exploit. Some data may be shortly unavailable on Wednesday.
  • NL-T1:
    • On Saturday morning around 05:17 the storage of one of our dCache pool nodes became inaccessible. This node contains ATLAS, Alice and LHCb data. This morning we managed to get the data back on line but at the moment we are still testing the system. We plan to put it back into production this afternoon. A downtime has been declared for this. The cause of this problem was most likely file system corruption due to a failed disk controller last week. The file system corruption has been repaired. Alexander was disconnected at the time of this report...
  • NRC-KI: not connected
  • OSG: Problem with site Purdue. It is a decommissioned resource in OIM but it still shows in SiteDB. The site admins can't update CMS web tools. Christoph said they should obtain the status of "executive" to be able to do this.
  • PIC: not connected
  • RAL: Two downtimes are registered in GOCDB for patches and tape robot intervention.
  • TRIUMF: NTR

  • Most CERN services:
    • Major datacentre network upgrade planned for November 3rd 06.00 - 08.30 - NOTE the date and time change! (to avoid major intervention on electrical supply to the site on the 2nd).
      • Expect a 3 to 7 minute outage on many services in the CERN CC during the indicated period as routers are rebooted.
      • Many batch / data nodes, myproxy, voms and other Grid services, together with multiple CERN site services affected. Many service VMs affected (owners have been notified).
      • Ref ITSSB: OTG:0033256
      • Ref GOCDB: https://goc.egi.eu/portal/index.php?Page_Type=Downtime&id=21672
  • CERN computing services:
    • VOMS intervention to upgrade service. Planned downtime Tuesday 8th November 10.00 - 12.00 for upgrade.
  • CERN storage services:
    • FTS will suffer a 3 to 7 minutes loss of network connection once or twice during the interval window of a Data Center network intervention Thu 3 Nov from 6 to 8:30 CET (OTG:0033736). Clients connections and running transfers will be affected.
    • FTS upgrade to v 3.5.7 (OTG:0033851), Tue 8 Nov 9:30-10:30.The upgrade will be transparent to clients
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • NTR. Apologies that we cannot be present today.
  • MW Officer: NTR
  • Networks: absent
  • Security: All new kernels for CVE-2016-5195 are out. Recommendation: schedule kernel updates & reboot WNs. Install mitigation in the mean time. EGI deadline (for update OR mitigation): today.
AOB:

NB! MW Readiness WG meeting Wed @ 4pm CET and Ops Coord Thu @ 3.30 pm! All agendas from index http://indico.cern.ch/category/4372/

Edit | Attach | Watch | Print version | History: r25 < r24 < r23 < r22 < r21 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r25 - 2016-11-02 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback